OpenStudy (anonymous):

Solve for x. 15-8x=1

jimthompson5910 (jim_thompson5910):

We have \[\Large 15-8x=1\] what's the first step in solving for x?

jimthompson5910 (jim_thompson5910):

1-15 is not -16 sarah

OpenStudy (anonymous):

sarah you are wrong

OpenStudy (anonymous):

first we subtract 15 on both sides

jimthompson5910 (jim_thompson5910):

that's only if it were -1-15

jimthompson5910 (jim_thompson5910):

good, so \[\Large 15-8x=1\] \[\Large 15-8x-15=1-15\] \[\Large -8x=-14\]

jimthompson5910 (jim_thompson5910):

then what?

OpenStudy (anonymous):